OpenStudy (anonymous):

What is the length of its minor axis? (x=5)^2 + (X+1)^2 ____________________ 144 225

OpenStudy (amistre64):

minor means small, so take the sqrt of the smallest denominator for half the length of the minor axis

OpenStudy (imstuck):

This is an ellipse, and the basic formula is\[\frac{ (x-h)^{2} }{ a ^{2} }+\frac{ (y-k)^{2} }{ b ^{2} }=1\]or\[\frac{ (x-h)^{2} }{ b ^{2} }+\frac{ (y-k)^{2} }{ a ^{2} }=1\]depending upon which is your major axis. The a is always ALWAYS greater than b, so whichever axis the a is under is your major axis. Having said that, the a is under the x term, so the x axis is your major axis, the y is your minor. Take the square root of b^2 (144) to get the length of the minor axis. That would be 12, then double it for the full length.